OpenStudy (anonymous):

isosceles trapezoid with large base=11 and small base = 5. calculate height

OpenStudy (callisto):

Is the area given?

OpenStudy (maheshmeghwal9):

The length of each diagonal is, according to Ptolemy's theorem, given by \[ p=\sqrt{ab+c^2}\] where a and b are the lengths of the parallel sides AD and BC, and c is the length of each leg AB and CD

OpenStudy (maheshmeghwal9):

Therefore; The height is, according to the Pythagorean theorem, given by \[ h=\sqrt{p^2-\left(\frac{a+b}{2}\right)^2}=\tfrac{1}{2}\sqrt{4c^2-(a-b)^2}.\]

OpenStudy (maheshmeghwal9):

It is copy paste from wiki I swear lol http://en.wikipedia.org/wiki/Isosceles_trapezoid

OpenStudy (maheshmeghwal9):

but i didn't understand the theory. can u plz solve the question @Callisto :)

OpenStudy (callisto):

@maheshmeghwal9 To be practical, we are given length of two sides only. How to find the third side?

OpenStudy (maheshmeghwal9):

ya I am too confused @ that

OpenStudy (anonymous):

neparalels sides are not given. All I know are large base = 11 and small base = 5.

OpenStudy (callisto):

Any diagram / other information?

OpenStudy (lgbasallote):

well it says isosceles :)

OpenStudy (anonymous):

Unfortunately no...Is a issoceles trapezoid

OpenStudy (maheshmeghwal9):

But without the third side or area How can we find height?

OpenStudy (lgbasallote):



OpenStudy (lgbasallote):

maybe that figure can help?

OpenStudy (anonymous):